The guitar string wire in item 2 was 1.60 m long with a diameter of 0.20 cm. Approximately what tension force would break it?


1
Expert's answer
2021-11-07T19:25:46-0500
"F=A\sigma=A\epsilon E=A\frac{\Delta L}{L}E=\frac{\pi d^2}4·\frac{\Delta L}{L}·E,\\\space\\
F=1.96·10^{-6}\Delta L·E."

Substitute the values of E and change in length from item 2 to get the answer.